silvernight269:

a local newspaper charges by the number of words for advertising. write a function to represent the advertising costs 10 words:7.50, 15 words:8.75 20 words:10.00, 25 words:11.25

Vocaloid:

hint: what do you think your independent and dependent variables are?

silvernight269:

is this still for the Arithmetic sequence?

Vocaloid:

well I guess you could treat this as a sequence but you would probably want to write it as a function instead

Vocaloid:

like f(x) = __________

silvernight269:

then i wouldn't know I'm still in middle school and i just started this unit

Vocaloid:

when we have one "variable" (in this case, the number of words) that determines the other (the price), we can write "function" which is a rule that tells us what the value of the independent variable is, if we input a certain dependent variable

Vocaloid:

(I realize that is kind of confusing so let me give an example)

Vocaloid:

Let's say a store is selling candy for 10 cents a piece. There are two things that are "varying" here; the number of candies, and the total amount paid.

Vocaloid:

Just using basic arithmetic (multiplication) we can determine the money paid simply by multiplying the price per candy by the number of candies so 1 candy --> 0.10 * 1 = $0.10 or 10 cents 2 candy -- > 0.1 * 2 = $0.20 or 20 cents

Vocaloid:

now, we want to figure out a "function" which will tell us the price paid for a generic number of candies, x

silvernight269:

oooohhh i get it

Vocaloid:

we can say: price = 0.10 * number of candies or, in more "mathematical" terms f(x) = 0.10x make sense so far?

silvernight269:

yes

Vocaloid:

so if we let x = 1 candy we get f(1) = 0.10*1 = $0.10 which is the same as our manual calculations from earlier

silvernight269:

yes

Vocaloid:

Now, your problem is a bit more difficult but the same principle applies

silvernight269:

so this is what i got so far a(n)=5.25+(n-1)0.25

silvernight269:

i think this is right but not 100% sure

Vocaloid:

If your teacher wants you to use series/sequence notation, then sure, we can go with that

silvernight269:

how would you do it?

Vocaloid:

yeah, your math looks right to me, good job ^^

Vocaloid:

hm

Vocaloid:

Doing the math I end up with f(x) = 5 + 0.25x But if you haven't started learning function notation, then go with your original answer.

silvernight269:

i did but i learned it this way too

silvernight269:

but I'm wondering if there is a more easier way to do these kind of problems

Vocaloid:

have you learned anything about slope-intercept form yet?

silvernight269:

yes

Vocaloid:

if you treat: 10 words:7.50, 15 words:8.75 as two points (10,7.50) and (15,8.75) you can just calculate the slope and y-intercept between them

silvernight269:

i see

Vocaloid:

so using our slope formula m = (y2-y1)/(x2-x1) we get (8.75-7.50)/(15-10) = 0.25

silvernight269:

yes

Vocaloid:

and then you just substitute that m into f(x) = mx + b to get your y-intercept

silvernight269:

oh

Vocaloid:

(keep in mind this only works for linear equations - or sequences with a common difference)

silvernight269:

ok. thank you so much!

Vocaloid:

np